A bicycle store costs $2400 per month to operate. The store pays an average of $60 per bike. The average selling price of each bike is $120. How many bicycles must the store sell?
(please show a system of equations)

Answer by ikleyn(52921)

.

Your equation is THIS

    2400 + 60n = 120n.


Its solution is


    2400 = 120n - 60n

    2400 = 60n

    n    = 2400/60 = 40.


ANSWER.  40 bicycles must be sold.

-------------

Solved.

That's all.

No system of equation.

You get the FULL solution, in short and clear form.

In shortest and clearest form.

Have fun and be happy (!)